Oil flow in a journal bearing can be treated as parallel flow between two large isothermal plates with the lower plate moving at a constant velocity of V =12 m/s and the other stationery. Consider such a flow with a uniform spacing of d=0.5mm between the plates. The temperatures of the upper and lower plates are T1=50 °C and T2=15°C, respectively. Use thermal conductivity k=0.145 W/m.K and dynamic viscosity µ =0.8374 kg/m.s for the oil.

(a) Simplify the mass continuity, momentum, and energy equations. (5 points)

(b) Determine the velocity and temperature distribution in the oil. (5 points)

(c) Find the maximum temperature and where it occurs. (5 points)
Answer #1

Average temperature at bearing as Te Title Ti= 80445 2 To= 32.5 given Ka) 8.145 wink o dx 2 20 dx Miss 0.8374 Ns/m² ConsiderOn integration Од – С. ५ argain integrate (9) - G9 +12 Екофа» Now to) -o 4(L) y clearance of beely O=Cҳо +12 2 -0 V z Cıxl to(32= To-liter? ОТ ay integration MY Z + c3 again integration & Tly=-uy ²1 %) + (zyrly- boundary conditions I (0) = Ti 2K T(L)7 t & F. 2 272 키 료? a-ti)y E-)y tury y² ) 시? ㅠ 쐞는 + 쟈 + (27) J velocity distribution is July) - 보 temperature distribution isso location of tmav 0.27 mm ya (max temps 2 +TI (I-IOy+ le2 Tmax ₂ 3 ) (분 Vy T 2k Y= 0.271 0.000271 7 to OaQba27) 0.0005 0.83